A dice is rolled two times. Find the probability of getting a composite number on first roll and a prime number on second roll.
Show answer & explanation
Correct answer: B
Explanation
The correct answer is option 2 i.e. 1/6
P(Getting a composite number) = 2/6 = 1/3
P(Getting a prime number) = 3/6 = 1/2
P(Getting a composite number on first roll and a prime number on second roll) = 1/3 × 1/2 = 1/6